Peter received commissions of $150 and $225. How much be receive from a third sale so his total commission will exceed $500?
Answer by doukungfoo(195)

To find the total amount of commissions earn you need to add them all together.
In this problem Peter has already received two commissions which are $150 and $225. The unknown amount is the third commission. Lets set this amount equal to x. Now we can write an equation to find the unknown amount x.
$150 + $225 + x = $501
first add the two known commissions together
$375 + x = $501
now to solve for x subtract $375 from both sides of the equal sign
$375 - $375 + x = $501 - $375
x = $126
So thats it. The third commission must be at least $126 for the total commissions earned to exceed $500